Just solved (Task 2)

Just solved (Task 2)

You work as a technical engineer in a company that designs and installs sound systeins, such as speakers for large halls and3. Use Excel or MATLAB to draw the function v and its first and second derivatives v and v, for 0<x<3.5. Then use these curTask 2: A A large speaker that you are planning to use, has a huge diaphragm connected to a spring, the force required to com

Just solved (Task 2)
Answer #1

TA w=ffde = f(covox+ sox3) de 5002? + 12.521+4 at a=0.tm at x=0, W=0 W = Sob(0,01) +1215X10 4 » G=0 = 5+ 0.00125 = 5.00025 Jowith Hence conside the variation. te viscosity of a lubricant decreases increase in Temperature idering I (06)] HaTTBT в Aeri